A long, straight wire of diameter 5.0 mm carries a uniformly distributed current of
15 A. At what distance from the axis of the wire will the magnitude of
B

be
maximum? Justify your answer
1
Expert's answer
2020-03-04T10:01:48-0500
"B=\frac{\mu_0I}{2\pi x},x\geq r"

Maximum will be at the x=r:


"B=\frac{4\pi \cdot 10^{-7}(15)}{2\pi (0.0025)}=0.0012\ T=1.2\ mT"